OpenStudy (czesc):

Determine the area of the inner loop of the polar curve r = 2 + 4 cos (θ).

OpenStudy (irishboy123):

\(\frac{1}{2}\int\limits_{2 \pi / 3}^{4 \pi /3} (2 + 4 \cos \theta ) ^2 \; d \theta\) and you can i think, maybe, half angle the integrand to make it a smidge easier